OpenStudy (anonymous):

FAN AND MEDEAL Find the product.Simplify 1. 1/10 x 3/4 A 15/2 B 2/15 C 3/40 D 40/3 2. 4/5 OF 40 A 32 B 1/32 C 12 D 44/5 3. 1/3 OF 3 A 1/18 B 1/3 C 1 D 1/9 4. Betty had 1/7 yard of ribbon. She used 1/6 of it. How many yards of ribbon did she use? 5. Suppose you have 15 muffins. How many muffins are left after you give a friend 1/3 of them? A 12 muffins B 10 muffins C 45 muffins D 5 muffins

myininaya (myininaya):

\[\frac{a}{b} \times \frac{c}{d}=\frac{ a \cdot c}{b \cdot d}\]

myininaya (myininaya):

can you try the first one knowing this?